• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name ransome

    Log entry time 00:03:11 on June14,2000

    Entry number 43671

    keyword=shift summary

    16:00 Ed Folts et al. still in the hall working on Q3
    20:20 Ed reports they will close up and we can try again
    Q2 and Q3 turned on, run up to 1600 A, set back to P=3.66
    Dipole started to go down, but we were able to turn it back on.
    20:45 carbon target in, waiting for beam
    problems with orbit, MCC goes to pulsed beam
    21:45 start carbon pointing
    22:25 moved to LH2 15 cm, requested higher beam
    we then realized we forgot to to 15 cm dummy. Decided to
    just go on for now.
    MCC has problems getting beam back, problem with raster
    23:20 finally beam back, Eugene comes in, we decide to
    proceed with Moller
    23:35 Q3 run down to 0 amps, turned off Voltmeters reading 62.6 and 61.5
    Moller in progress
    00:02 Voltmeters increasing,now 63.3, 63.6

    Run
    2012 Carbon, raster off 5uA test if things are running
    2013 Harp scan, carbon target
    2014 5 minute Carbon run raster off, 5 uA
    2015 4 cm dummy, raster off 5 uA
    2016 10 cm dummy, raster off 5 uA